Shift report 20020123 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 23 2002 - 07:50:20 EST

  • Next message: Flemming Videbaek: "end of pp running"

    ------------------------------------------------------------
    Date: Wed Jan 23 07:50:20 2002, Shift: 00:00-08:00
    Supervisor: hiro, with: 
    Log book#12 pages 106 - 107
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Bean until 04:00 
        At 4, consequtive pol measurement followed by
        spin flipper test.
        At 5:45 Yellow QLI, TV indicates that no beam
        until 9:30
    
    2.  run 6740, 6741 
        (MRS 90 deg 6k B field, FS 12 1/2 A field as I was
         told to keep this setting from the previous shift)
    
    3.  NONE noticed.
    
    Only a few more days (or none???) left... (Is the last day
    set aside for pp2pp?)
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Jan 23 2002 - 07:51:00 EST